OpenStudy (anonymous):

Allie currently has an account balance of $2,265.96. She opened the account 13 years ago with a deposit of $1,227.43. If the interest compounds daily, what is the interest rate on the account?

OpenStudy (hunus):

The formula for compounded interest is \[A=P(1+\frac{r}{n})^{nt}\] Where n is the number of times per year it is compounded So the equation you will need to use is \[A=P(1+\frac{r}{365})^{365t}\]

OpenStudy (anonymous):

i got 1.9%

OpenStudy (hunus):

I got something different. Will you show me your steps?

OpenStudy (anonymous):

i got 4.7%

OpenStudy (anonymous):

2265.96(1+r/365)365*12

OpenStudy (anonymous):

...........

OpenStudy (hunus):

Okay is it 12 or 13 years?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

intrest rate

OpenStudy (hunus):

no? In your question you put 13 in your step you put 12

OpenStudy (hunus):

I know. But I need to know if t = 12 or t = 13

OpenStudy (anonymous):

whats r

OpenStudy (hunus):

\[2265.96=1227.43(1+\frac{r}{365})^{365*13}\] \[\frac{2265.96}{1227.43}=(1+\frac{r}{365})^{365*13}\] \[\log(1.8461)=\log((1+\frac{r}{365})^{365*13})\] \[\log(1.8461)=365*13*\log((1+\frac{r}{365}))\] \[\frac{\log(1.8461)}{365*13}=1.000129=\log(1+\frac{r}{365})\] \[10^{1.000129}=1+\frac{r}{365}\] \[10^{1.000129}-1=0.000129=\frac{r}{365}\] \[r=0000129*365=0.0471\]